Please help me solve.
2x^2-4x+16=0

Answer by checkley71(8403)

2X^2-4X+16=0
2(X^2-2X+8)=0
2(X-4)(X+2)=0
X-4=0
X=4 ANSWER.
X=2=0
X=-2 ANSWER.